# Exploring the extent of similarities in software failures across industries using LLMs

### v1

@@ Abstract
Background/Motivation:
The rapid evolution of software development necessitates enhanced safety measures. Extracting information about software failures from companies is becoming increasingly more available through news articles.
Problem Statement:
This research utilizes the Failure Analysis Investigation with LLMs (FAIL) model to extract industry-specific information. Although the FAIL model’s database is rich in information, it could benefit from further categorization and industry-specific insights to further assist software engineers.
Methods:
In previous work news articles were collected from reputable sources and categorized by incidents inside a database. Prompt engineering and Large Language Models (LLMs) were then applied to extract relevant information regarding the software failure. This research extends these methods by categorizing articles into specific domains and types of software failures. The results are visually represented through graphs.
Results:
The analysis shows that throughout the database some software failures occur significantly more often in specific industries. This categorization provides a valuable resource for software engineers and companies to identify and address common failures.
Conclusion:
This research highlights the synergy between software engineering and Large Language Models (LLMs) to automate and enhance the analysis of software failures. By transforming data from the database into an industry specific model, we provide a valuable resource that can be used to identify common vulnerabilities, predict potential risks, and implement proactive measures for preventing software failures. Leveraging the power of the current FAIL database and data visualization, we aim to provide an avenue for safer and more secure software in the future.
@@ Introduction
The Securities and Exchange Commission (SEC) recently made a ruling requiring public companies to disclose cyber-security related failures. This means that any breach in security deemed to have potential significant impact or harm must be reported within 4 business days of the incident. Public companies must also create reports annually of cyber-security risk management strategies, and governance [7]. With this increased transparency requirement in the industry, the FAIL model, and industry specific insights will only become more useful and prevalent in the future.
@@
We began by creating a prompt from which we collect data from the database, and categorize the data into industry-specific failures. The initial taxonomy was adopted from [8, 9, 10] which give taxonomies for categorizing software failures. The temperature of the LLM that we used was manually set to 0 to increase consistency, and we used chatgpt-3.5-turbo. Initially the prompt contained general instructions, such as ”categorize these prompts by type of software failure.” However the prompt’s performance was sub optimal. Effective prompts need to be clear and concise to guide the Large Language Model (LLM) in producing the desired output [2]. The initial prompt generated results with high variability or ”randomness”. To improve accuracy, we revised the prompt with a specific list of common software failures manually curated based on descriptions in the database, and taxonomies widely used in the industry for categorizing software failures. [10, 9] We ensured the prompt only returned the type of software failure. This adjustment yielded more consistent results by the LLM, but our approach was still producing some incorrect categorizations. After refining the prompt and using few-shot prompting, the prompt performed significantly better. Through the use of a confusion matrix and a small dataset of 90 articles, the overall average accuracy of the prompt was 76 percent (reference Figure 1 for an example of the matrix used). With this accuracy we let the script run through the database, and categorize each software failure by industry. Results are shown visually through bar graphs.
@@ Acknowledgements
NSF CPS Frontier: Cognitive Autonomy for Human CPS: for funding this project.

### v2

@@ Abstract
Background/Motivation: The rapid evolution of software development necessitates enhanced safety measures. Extracting information about software failures from companies is becoming increasingly more available through news articles.
Problem Statement: This research utilizes the Failure Analysis Investigation with LLMs (FAIL) model to extract industry-specific information. Although the FAIL model’s database is rich in information, it could benefit from further categorization and industry-specific insights to further assist software engineers.
Methods: In previous work news articles were collected from reputable sources and categorized by incidents inside a database. Prompt engineering and Large Language Models (LLMs) were then applied to extract relevant information regarding the software failure. This research extends these methods by categorizing articles into specific domains and types of software failures. The results are visually represented through graphs.
Results: The analysis shows that throughout the database some software failures occur significantly more often in specific industries. This categorization provides a valuable resource for software engineers and companies to identify and address common failures.
Conclusion: This research highlights the synergy between software engineering and Large Language Models (LLMs) to automate and enhance the analysis of software failures. By transforming data from the database into an industry specific model, we provide a valuable resource that can be used to identify common vulnerabilities, predict potential risks, and implement proactive measures for preventing software failures. Leveraging the power of the current FAIL database and data visualization, we aim to provide an avenue for safer and more secure software in the future.
@@ Introduction
The Securities and Exchange Commission (SEC) recently made a ruling requiring public companies to disclose cybersecurity-related failures. [7] This means that any breach in security deemed to have a significant impact or potential harm must be reported within four business days of the incident. Public companies must also create annual reports of cybersecurity risk management strategies and governance. With this increased transparency requirement in the industry, the FAIL model and industry-specific insights will only become more useful and prevalent in the future.
Previous research has laid the foundation for using Large Language Models (LLMs) to analyze software failures. In ”Can Large Language Models Analyze Software Failures in the News? An End-to-End Automated Pipeline with FAIL,” [1] Anandayuvaraj demonstrated the potential of LLMs to automate the process of analyzing news articles and extracting relevant information about software failures. This work shows the insurmountable importance of the use of LLMs in software failure analysis. With this study many different avenues open in terms of research such as why software failures happen, methods of prevention, and common software failure risks.
Another significant study, ”Reflecting on Recurring Failures in IoT Development,” [11], highlighted the challenges specific to Internet of Things (IoT) Development. Anandayuvaraj identified common patterns of failures in IoT systems and emphasized the importance of incorporating lessons learned into future designs to prevent recurring consistent issues. This research not only enforces the necessity of industry-specific analysis to address unique challenges, but also supports assumptions that some failures could be dominant in software development.
In ”Incorporating Failure Knowledge into Design Decisions for IoT Systems: A Controlled Experiment on Novices,” [12] Anandayuvaraj explored how failure knowledge could be integrated into the design process. This study focused on evaluating the influence of failure-related learning techniques. The study revealed that providing novice developers with insights into common failures can significantly enhance the quality of IoT systems. This experiment supports the idea that sharing industry-specific failure knowledge can improve software design and development.
In ”Reflections on Software Failure Analysis,” [14] Amusuo explored different software failure studies along with common shortcomings of these previous studies. They found that studies have a bias towards open source data, and there is a lack of consistency in taxonomies used in identifying failures. This study highlights the importance of a universal taxonomy for comparison throughout different studies, as well as a need for improved replicability amongst studies.
Finally, ”An Empirical Study on Using Large Language Models to Analyze Software Supply Chain Security Failures” [13] examined the applicability of LLMs in identifying vulnerabilities in the software supply chain. Anandayuvaraj’s findings highlighted the effectiveness, and some possible drawbacks of LLMs in identifying patterns and potential risks within supply chain security.
These works collectively illustrate the potential of leveraging LLMs and industry-specific insights in understanding and preventing software failures. By building upon this foundation, the current research aims to further refine these methodologies and enhance the utility of the FAIL model in delivering insights for software engineers and companies.
@@
We began by creating a prompt to collect data from the database and categorize it into industry-specific failures. The initial taxonomy was influenced by previous works [8, 9, 10], which provide frameworks for categorizing software failures. To increase consistency, we manually set the temperature of the LLM (chatgpt-3.5-turbo) to 0.
Initial Prompt Design
The initial prompt included general instructions such as:
You are an AI trained to identify the type of software failure based
on the Cause provided. Use the context in the cause to determine the
type of software failure
However, this prompt’s performance was suboptimal, as effective prompts need to be clear and concise to guide the Large Language Model (LLM) in producing the desired output [2]. The initial prompt generated results with high variability or ”randomness.”
Iteration 1: Specific Failure List
To improve accuracy, we revised the prompt with a specific list of common software failures, manually curated based on descriptions in the database and taxonomies widely used in the industry for categorizing software failures [10, 9]. The revised prompt was:
You are an AI trained to identify the type of software failure based on
the Cause provided. Use the context in the cause to determine the type
of software failure from the following list:
        - Data Breach
        - Functionality Bug
        - UI/UX Bug
        - Regression Bug
        - Outage
        - Security Vulnerability
        - Performance Issue
        - Integration Issue
        - Non-Software Cause
        - Other
Your task is to read the cause below and determine the type of software
failure from the given list. Only return the type of failure without any
additional text or explanation. If the cause does not fit any category
in the list, return "Other".
This adjustment yielded more consistent results, but some incorrect categorizations persisted.
Iteration 2: Few-Shot Prompting
To further refine the prompt, we incorporated few-shot prompting, providing examples of correct categorizations. The updated prompt was:
You are an AI trained to identify the type of software failure based on
the Cause provided. Use the context in the cause to determine the type
of software failure from the following list:
        - Data Breach
        - Functionality Bug
        - UI/UX Bug
        - Regression Bug
        - Outage
        - Security Vulnerability
        - Performance Issue
        - Integration Issue
        - Non-Software Cause
        - Other
Your task is to read the cause below and determine the type of software
failure from the given list. Only return the type of failure without any
additional text or explanation. If the cause does not fit any category
in the list, return "Other"
Examples:
Cause: The faulty computer software in the Chinook helicopter was described
as "positively dangerous" and had deficiencies that meant the pilot’s full
control of the engines "could not be assured".
Functionality Bug
Cause: Software bug causing Windows Mobile devices to incorrectly date
incoming SMS messages to 2016.
Regression Bug
Cause: Lack of Adobe Flash support on Apple’s Safari browser for iPad, iPhone,
and iPod Touch devices.
Non-Software Cause
This iteration significantly improved the prompt’s performance.
Evaluation and Results
We evaluated the prompt’s accuracy using a confusion matrix and a small dataset of 90 articles. The overall average accuracy of the prompt was 76 percent (reference Figure 1 for an example of the matrix used). With this accuracy, we ran the script through the database to categorize each software failure by industry. Results are shown visually through bar graphs. Explanation
@@ Acknowledgements
  • This project was funded by NSF award CNS-1836952 via the SIRI program
  • Peer: Bryan Zhang for useful feedback/reviews during the creation of this study